# MGMCpaper.bib

### v2

@@ 5.3 Grid independent convergence
As shown in Section 4, one of the key advantages of the MGMC algorithm is that it shows grid-independent convergence as $h_{L}\rightarrow 0$. This is in contrast to a naive Gibbs sampler, for which the integrated autocorrelation time and the root mean squared error at fixed sample size grow if the resolution increases. We now demonstrate this by computing several performance indicators of these two samplers. For all results in this section the prior is the FEM discretisation of the shifted Laplace operator $\mathcal{A}^{\mathrm{(SL)}}=-\Delta+\kappa^{2}I$ defined in (5.2) in $d=2$ dimensions.
@@ 5.3.1 Convergence to the target distribution
where the vector-space isomorphism $P_{L}:\mathbb{R}^{n_{L}}\rightarrow V_{L}$ is defined in Section 2.3 and $x_{\text{centre}}$ is the centre of the domain $\Omega$; the radius $R=0.025$ is the same as for the observations that define the posterior distribution (see (5.5)). Note that $F$ is a linear operator of the form given in (2.16) and as a consequence for each step $k$ in the Markov chain $z_{L}^{(m)}$ is a normal random variable with some mean $\mu_{L}^{(m)}$ and variance $(\sigma_{L}^{(m)})^{2}$. Under suitable conditions we have that $\mu_{L}^{(m)}\rightarrow\mu_{L}$ and $\sigma_{L}^{(m)}\rightarrow\sigma_{L}$ as $k\rightarrow\infty$ (see Theorem 4.9). To quantify the rate of convergence we consider the ratios
@@
$\widehat{\mu}_{L}^{(m)}$ $:=\frac{1}{n_{\text{samples}}}\sum_{j=1}^{n_{\text{samples}}}z_{L;j}^{(m)}\approx\mathbb{E}[z_{L}^{(m)}]=\mu_{L}^{(m)}\qquad\text{with $z_{L;j}^{(m)}=F(\theta_{L;j}^{(m)})$}$ (5.10)
$(\widehat{\sigma}_{L}^{(m)})^{2}$ $:=\frac{1}{n_{\text{samples}}-1}\sum_{j=1}^{n_{\text{samples}}}(z_{L;j}^{(m)}-\widehat{\mu}_{L}^{(m)})^{2}\approx\mathrm{Var}[z_{L}^{(m)}]=(\sigma_{L}^{(m)})^{2}$
Replacing $\mu_{L}^{(m)}\mapsto\widehat{\mu}_{L}^{(m)}$, $\sigma_{L}^{(m)}\mapsto\widehat{\sigma}_{L}^{(m)}$ in (5.9), we can compute the estimators $\widehat{R}_{L}^{(m)}$, $\widehat{Z}_{L}^{(m)}$ with associated statistical errors that arise from the finite sample size. Fig. 1 shows a plot of these estimators for the first 16 steps in the Markov chain. Note that for both samplers the variance converges faster than the mean, which is consistent with Theorem 4.10. For the Gibbs sampler, convergence is extremely slow. In contrast, for MultigridMC the estimated mean $\mu_{L}^{(m)}$ and covariance $(\sigma_{L}^{(m)})^{2}$ can not be distinguished from the mean $\mu_{L}$ and variance $\sigma_{L}^{2}$ of the target distribution within statistical errors after a small number of steps.
Figure 1: Convergence of the mean $\mu_{L}^{(m)}$ and variance $(\sigma_{L}^{(m)})^{2}$ for the Gibbs sampler (red dashed lines) and for the MultigridMC sampler (blue solid lines). The plot shows the estimators $\widehat{R}_{L}^{(m)}$ and $\widehat{Z}_{L}^{(m)}$ for the quantities defined in (5.9). The grid is of size $128\times 128$ and the number of independent Markov chains is $n_{\text{samples}}=100,000$.
@@
Figure 2: Dependence of the convergence rates $\widehat{\rho}_{L}$ (left) and $\widehat{\zeta}_{L}$ (right) defined in (5.11) on the resolution for different correlation lengths $\kappa^{-1}$. For $\kappa^{-1}=0.01$ the mean $\mu_{L}^{(m)}$ converged so rapidly for both samplers that the convergence rate $\widehat{\rho}_{L}$ could not be measured reliably with the given statistics.
@@ 5.3.2 Autocorrelations
While the Cholesky sampler produces inherently independent samples, the states $\theta_{L}^{(m)}$ in the Markov chains generated by the Gibbs sampler and the MGMC update in Alg. 1 are inherently correlated. To explore this, we discard the first $n_{\text{warmup}}=1000$ samples from the Markov chain to account for burn-in of the chain so that to a good approximation $\theta_{L}^{(0)}\sim\mathcal{N}(\widetilde{A}_{L}^{-1}f_{L},\widetilde{A}_{L}^{-1})$. This will result in a time series $z_{L}^{(0)},z_{L}^{(1)},z_{L}^{(2)},\dots$ with the observable $z_{L}^{(m)}$ defined as in (5.8). We compute the lagged autocorrelation function $\widehat{\Gamma}_{z}(t)/\widehat{\Gamma}_{z}(0)$ defined by
$\quad\widehat{\Gamma}_{z}(t)$ $:=\frac{1}{n_{\text{steps}}-t}\sum_{m=0}^{n_{\text{steps}-1-t}}(z_{L}^{(m)}-\widehat{z}_{L})(z_{L}^{(m+t)}-\widehat{z}_{L})$ (5.12)
@@
To further quantify autocorrelations in the Markov chain we estimate the integrated autocorrelation time (IACT) defined in (4.28). Computing the IACT is notoriously difficult if the autocorrelations are strong. Here we use the method in [wolff2004monte] to estimate the IACT, i.e. we compute the estimator
@@
where the optimal choice of window size $W$ is described in [wolff2004monte, Section 3.3]. As stated in Theorem 4.12, the IACT for MGMC is grid-independent. As in Fig. 2 we also explore the dependence of IACT on the correlation length $\kappa^{-1}$. Fig. 3 (right) shows the IACT for both the MGMC sampler and the standard Gibbs sampler as the resolution increases and for a range of different $\kappa^{-1}$.
@@ 6 Conclusion
In this paper we have presented a rigorous analysis of the Multigrid Monte Carlo approach originally proposed by Goodman and Sokal in [goodman1989multigrid]. We have extended the analysis to the important case of conditioned Gaussian random fields in a linear Bayesian setting. We have shown that the rate of convergence to the targest distribution and the integrated autocorrelation time is grid-independent. To achieve this, we relied on the inherent connection between solvers for sparse linear systems and samplers for multivariate normal distributions discussed in [Fox.C_Parker_2017_AcceleratedGibbsSampling]. This allowed us to bound convergence rates with tools from standard multigrid theory which, however, had to be extended to the Bayesian setup considered here. Our cost-analysis confirms that the cost for one MGMG update grows linearly in the number of unknowns. As a consequence, the algorithm is optimal in the sense that the cost for generating a single independent sample is also proportional to the number of unknowns.
The theoretical results are confirmed by a set of numerical experiments, which demonstrate that MGMC is robust and efficient as the resolution increases. On finer grids, MGMC is always significantly faster than a standard Gibbs sampler. In three dimensions and on larger lattices, it beats even the Cholesky method, which demonstrates that our approach is particularly promising in higher dimensions. We are able to sample very rough fields with precision operators $-\Delta+\kappa^{2}I$ and $(-\Delta+\kappa^{2}I)^{2}$ without problems, but find that MGMC is most efficient in the former case.
@@
There are several ways of extending the work in this paper. While we only considered relatively small problems here, significantly larger problems can be simulated with the parallel implementation which will be described in a subsequent publication. For simplicity, we have also limited ourselves to a linear setting where in particular the quantity of interest and the observation operators are bounded linear functionals. It would be interesting to consider more general cases with likelihoods given by $\propto\exp\left[-\frac{1}{2}(\mathcal{B}_{\text{NL}}(\theta)-\mu)^{\top}\Gamma^{-1}(\mathcal{B}_{\text{NL}}(\theta)-\mu)\right]$ for some non-linear $\mathcal{B}_{\text{NL}}$. Similarly, the generalisation to non-Gaussian priors could be considered. While the original MGMC algorithm in [goodman1989multigrid] is written down for the general non-Gaussian case and these cases should be covered in principle, in practice it will likely have to be adapted to be efficient.
@@ Acknowledgements
[…]
@@ A.2 Representation of the MGMC iteration
Multiplying the first sum in (A.6) from the left by $A$, using $BA=I-S$ and evaluating the telescoping sum we find
@@
Let $A\in\mathbb{R}^{n\times n}$ be symmetric and invertible. Let the matrices $M,N\in\mathbb{R}^{n\times n}$ be such that $A=M-N$ and that $M$ is invertible. Then, with $S:=M^{-1}N$,
@@
:=$\mathscr{M}_{\ell}:=\{(m_{\ell},m_{\ell+1},\dots,m_{L-1})\mid 0\leq m_{\ell^{\prime}}\leq\gamma_{\ell^{\prime}+1}-1\;\text{for all $\ell^{\prime}=\ell,\dots,L$}\}$ (A.9)
@@
$W_{\ell}(m_{\ell},\dots,m_{L})$ $:=(S^{\mathrm{post}}_{\ell})^{\nu_{2}}\Bigl{(}I_{\ell-1}^{\ell}\sum_{m=1}^{\gamma_{\ell}}X_{\ell-1}^{m-1}W_{\ell-1}(\gamma_{\ell}-m,m_{\ell},\dots,m_{L})$ (A.12)
@@
$\qquad\qquad+\sum_{j=1}^{\nu_{2}}(S_{\ell}^{\mathrm{post}})^{j-1}(M_{\ell}^{\mathrm{post}})^{-1}w_{\ell}^{\mathrm{post}(\nu_{2}-j)}(m_{\ell},\dots,m_{L}),\qquad\text{$\ell\geq 1$}.$
@@ A.4 Equivalence of Multigrid and MGMC
Suppose (i) holds. Then, from [Young.D.M_book_1971, Theorem 3.5.1], the spectral radius of $X_{L}$ is less than $1$, so that $\lim_{m\to\infty}X_{L}^{m}=0$. But from (4.26a) and (4.26b) in Lemma 4.8, we have
@@
To show ((ii)$\implies$(i)) we show the contraposition. To show (ii) does not hold, expecting a contradiction suppose it does. Then, $\theta_{L}^{(m)}$ converges in distribution to $\mathcal{N}(A_{L}^{-1}f_{L},A_{L}^{-1})$ whatever the initial state is. Then, we must have $\lim_{m\to\infty}\mathbb{E}[\theta_{L}^{(m)}]=A_{L}^{-1}f_{L}$, as we will show below. But since (i) does not hold, we can choose $\theta_{L}^{(0)}$ such that $\lim_{m\to\infty}\mathbb{E}[\theta_{L}^{(m)}]\neq A_{L}^{-1}f_{L}$ because upon taking the expectation the iteration (A.30) is identical to (A.29), a contradiction.
@@ A.5 Convergence of distributions
see [murphy2022probabilistic, (6.32)]. Here, we note that for sufficiently large $m$, the continuity of the determinant together with Theorem 4.10 implies $\mathrm{det}(\Sigma_{m})\neq 0$. The continuity of the trace function and determinant implies $\lim_{m\to\infty}\operatorname{tr}(A_{L}\Sigma_{m})=n_{L}$ and $\lim_{m\to\infty}\frac{\mathrm{det}(A_{L}^{-1})}{\mathrm{det}(\Sigma_{m})}=1$. Hence,
@@
Similarly, $\lim_{m\to\infty}{D_{\text{KL}}\bigl{(}(A_{L}^{-1}f_{L},A_{L}^{-1})\parallel(\text{$\mu_{m}$, $\Sigma_{m}$})\bigr{)}=0}$ holds, where we again note $\mathrm{det}(\Sigma_{m})\neq 0$ for sufficiently large $m$. In view of the Pinsker’s inequality, the total variation distance converges to 00 as well, which implies the weak convergence. The weak convergence can be checked directly with the convergence of characteristic function. ∎

### v3

@@ 5.3 Grid independent convergence
As shown in Section 4, one of the key advantages of the MGMC algorithm is that it shows grid-independent convergence as $h_{L}\rightarrow 0$. This is in contrast to a naive Gibbs sampler, for which the integrated autocorrelation time and the root mean squared error at fixed sample size grow if the resolution increases. We now demonstrate this by computing several performance indicators of these two samplers. For all results in this section the prior is the FE discretisation of the shifted Laplace operator $\mathcal{A}^{\mathrm{(SL)}}=-\Delta+\kappa^{2}I$ defined in (5.2) in $d=2$ dimensions.
@@ 5.3.1 Convergence to the target distribution
where the vector-space isomorphism $P_{L}:\mathbb{R}^{n_{L}}\rightarrow V_{L}$ is defined in Section 2.3 and $x_{\text{centre}}$ is the centre of the domain $D$; the radius $R=0.025$ is the same as for the observations that define the posterior distribution (see (5.5)). Note that $F$ is a linear operator of the form given in (2.16) and as a consequence for each step $k$ in the Markov chain $z_{L}^{(m)}$ is a normal random variable with some mean $\mu_{L}^{(m)}$ and variance $(\sigma_{L}^{(m)})^{2}$. Under suitable conditions we have that $\mu_{L}^{(m)}\rightarrow\mu_{L}$ and $\sigma_{L}^{(m)}\rightarrow\sigma_{L}$ as $k\rightarrow\infty$ (see Theorem 4.11). To quantify the rate of convergence we consider the ratios
@@
$\widehat{\mu}_{L}^{(m)}$ $=\frac{1}{n_{\text{samples}}}\sum_{j=1}^{n_{\text{samples}}}z_{L;j}^{(m)}\approx\mathbb{E}[z_{L}^{(m)}]=\mu_{L}^{(m)}\qquad\text{with $z_{L;j}^{(m)}=F(\theta_{L;j}^{(m)})$}$ (5.10)
$(\widehat{\sigma}_{L}^{(m)})^{2}$ $=\frac{1}{n_{\text{samples}}-1}\sum_{j=1}^{n_{\text{samples}}}(z_{L;j}^{(m)}-\widehat{\mu}_{L}^{(m)})^{2}\approx\mathrm{Var}[z_{L}^{(m)}]=(\sigma_{L}^{(m)})^{2}$
Replacing $\mu_{L}^{(m)}\mapsto\widehat{\mu}_{L}^{(m)}$, $\sigma_{L}^{(m)}\mapsto\widehat{\sigma}_{L}^{(m)}$ in (5.9), we can compute the estimators $\widehat{R}_{L}^{(m)}$, $\widehat{Z}_{L}^{(m)}$ with associated statistical errors that arise from the finite sample size. Fig. 1 shows a plot of these estimators for the first 16 steps in the Markov chain. Note that for both samplers the variance converges faster than the mean, which is consistent with Theorem 4.12. For the Gibbs sampler, convergence is extremely slow. In contrast, for MGMC the estimated mean $\mu_{L}^{(m)}$ and covariance $(\sigma_{L}^{(m)})^{2}$ can not be distinguished from the mean $\mu_{L}$ and variance $\sigma_{L}^{2}$ of the target distribution within statistical errors after a small number of steps.
Figure 1: Convergence of the mean $\mu_{L}^{(m)}$ and variance $(\sigma_{L}^{(m)})^{2}$ for the Gibbs sampler (crosses) and for the MGMC sampler (filled circles). The plot shows the estimators $\widehat{R}_{L}^{(m)}$ (blue solid lines) and $\widehat{Z}_{L}^{(m)}$ (red dashed lines) for the quantities defined in (5.9). The grid is of size $128\times 128$ and the number of independent Markov chains is $n_{\text{samples}}=100,000$.
@@
Figure 2: Dependence of the convergence rates $\widehat{\rho}_{L}$ (left) and $\widehat{\zeta}_{L}$ (right) defined in (5.11) on the resolution for different correlation lengths $\kappa^{-1}$. For $\kappa^{-1}=0.01$ the mean $\mu_{L}^{(m)}$ converged so rapidly for both samplers that the convergence rate $\widehat{\rho}_{L}$ could not be measured reliably with the given statistics. While one would expect to see $2\times 3$ different curves for $\widehat{\rho}_{L}$ and $2\times 4$ curves for $\widehat{\zeta}_{L}$, the results for $\kappa=1.0$ and $\kappa=10.0$ lie on top of each other in the left figure and they almost overlap in the right figure, in particular for the MGMC sampler.
@@
We conclude this section by observing that although in exact arithmetic the Cholesky sampler produces states from the target distribution, in finite precision floating point arithmetic the ill-conditioning of the precision matrix leads to the accumulation of rounding errors. This implies that the generated samples are drawn from the wrong distribution for very large problems. However, for the problem sizes considered in our numerical experiments we did not find this to be an issue.
@@ 5.3.2 Autocorrelations
While the Cholesky sampler produces inherently independent samples, the states $\theta_{L}^{(m)}$ in the Markov chains generated by the Gibbs sampler and the MGMC update in Alg. 1 are inherently correlated. To explore this, we discard the first $n_{\text{burn-in}}=1000$ samples from the Markov chain to account for burn-in of the chain so that to a good approximation $\theta_{L}^{(0)}\sim\mathcal{N}(\widetilde{A}_{L}^{-1}f_{L},\widetilde{A}_{L}^{-1})$. This will result in a time series $z_{L}^{(0)},z_{L}^{(1)},z_{L}^{(2)},\dots$ with the observable $z_{L}^{(m)}$ defined as in (5.8). We compute the lagged autocorrelation function $\widehat{\Gamma}_{z}(t)/\widehat{\Gamma}_{z}(0)$ defined by
$\quad\widehat{\Gamma}_{z}(t)$ $=\frac{1}{n_{\text{steps}}-t}\sum_{m=0}^{n_{\text{steps}-1-t}}(z_{L}^{(m)}-\widehat{z}_{L})(z_{L}^{(m+t)}-\widehat{z}_{L})$ (5.12)
@@
To further quantify autocorrelations in the Markov chain we estimate the integrated autocorrelation time (IACT) defined in (4.28). Computing the IACT is notoriously difficult if the autocorrelations are strong. Here we use the method in [undefaax] to estimate the IACT, i.e. we compute the estimator
@@
where the optimal choice of window size $W$ is described in [undefaax, Section 3.3]. As stated in Theorem 4.14, the IACT for MGMC is grid-independent. As in Fig. 2 we also explore the dependence of IACT on the correlation length $\kappa^{-1}$. Fig. 3 (right) shows the IACT for both the MGMC sampler and the standard Gibbs sampler as the resolution increases and for a range of different $\kappa^{-1}$.
@@ 6 Conclusion
In this paper we have presented a rigorous analysis of the Multigrid Monte Carlo approach originally proposed by Goodman and Sokal in [undef]. We have extended the analysis to the important case of conditioned Gaussian random fields in a linear Bayesian setting. We have shown that the rate of convergence to the targest distribution and the integrated autocorrelation time is grid-independent. To achieve this, we relied on the inherent connection between solvers for sparse linear systems and samplers for multivariate normal distributions discussed in [undefae]. This allowed us to bound convergence rates with tools from standard multigrid theory which, however, had to be extended to the Bayesian setup considered here. Our cost-analysis confirms that the cost for one MGMG update grows linearly in the number of unknowns. As a consequence, the algorithm is optimal in the sense that the cost for generating a single independent sample is also proportional to the number of unknowns.
The theoretical results are confirmed by a set of numerical experiments, which demonstrate that MGMC is robust and efficient as the resolution increases. On finer grids, MGMC is always significantly faster than a standard Gibbs sampler. In three dimensions and on larger lattices, it beats even the Cholesky method, which demonstrates that our approach is particularly promising in higher dimensions. We are able to sample very rough fields with precision operators $\Delta+\kappa^{2}I$ and $(-\Delta+\kappa^{2}I)^{2}$ without problems, but find that MGMC is most efficient in the former case.
As stated above, the MGMC method is applicable in more general setups such as irregular domains or non-stationary distributions where other, potentially more efficient samplers based on spectral methods are not available.
@@
There are several ways of extending the work in this paper. While we only considered relatively small problems here, significantly larger problems can be simulated with the parallel implementation which will be described in a subsequent publication. For simplicity, we have also limited ourselves to a linear setting where in particular the quantity of interest and the observation operators are bounded linear functionals. It would be interesting to consider more general cases with likelihoods given by $\propto\exp\left[-\frac{1}{2}(\mathcal{B}_{\text{NL}}(\theta)-\mu)^{\top}\Gamma^{-1}(\mathcal{B}_{\text{NL}}(\theta)-\mu)\right]$ for some non-linear $\mathcal{B}_{\text{NL}}$. Similarly, the generalisation to non-Gaussian priors could be considered. While the original MGMC algorithm in [undef] is written down for the general non-Gaussian case and these cases should be covered in principle, in practice it will likely have to be adapted to be efficient.
The numerical results presented here were obtained for covariances of the form $(-\Delta+\kappa^{2}I)^{-\alpha}$ with $\alpha=1,2$. It would be interesting to also consider the case of fractional $\alpha$ since this corresponds to a more general class of Matérn covariance functions. However, it should be stressed here that further work is required to understand how MGMC can be applied to this setup.
While in this work we assumed that the number of observations $\beta$ is fixed and much smaller than the problem size $n_{L}$, it will be interesting to also consider the case where $\beta$ is large and grows with some power $n_{L}$. In this case the performance of the different samplers will likely depend on the distribution of the observations over the domain, and we expect MGMC to be particularly efficient if parts of the domain are not informed by the data. Note that if $\beta\propto n_{L}^{p}$ with $p>1$ the implementation needs to be adjusted to achieve optimal performance: for example, in this setup it might be advantageous to pre-compute the $n_{\ell}\times n_{\ell}$ matrix product $G_{\ell}B_{\ell}^{\top}$ in (3.15) instead of applying the $n_{\ell}\times\beta$ and $\beta\times n_{\ell}$ matrices $B_{\ell}^{\top}$ and $G_{\ell}$ in sequence. This will change the bounds on the computational cost and storage requirements in Section 3.6.
While here we considered operators that arise from the discretisation of a differential operator on a grid, statistical distributions can also be formulated on graphs. In this case, the precision operator has to be treated as a purely algebraic object with no underlying mesh hierarchy. However, the MGMC philosophy is still applicable in this case: as in algebraic multigrid (AMG) methods [undefaay, undefaaz], a hierarchy of coarse level spaces can be constructed in a purely algebraic way. The resulting sampler is obtained by replacing the AMG smoother by a suitable sampler. By following the same steps as in Section 4, the convergence theory of the sampler can be reduced to the well-established convergence theory of AMG. This also allows the easy application of MGMC on more complex domains and unstructured grids. A preliminary exploration of this idea, including some numerical experiments and parallel scalability studies, can be found in [undefaf].
@@ Acknowledgements
This work was funded as Exploratory Project 4.6 by the Deutsche Forschungsgemeinschaft (German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). We are grateful to Colin Fox for useful discussions, and thank Nils Friess for his careful reading of the manuscript and detailed comments. We would like to thank Jonathan Goodman for sending us many extremely helpful comments on the original submission, which have helped to improve this paper. This work made use of the nimbus cloud computer, and the authors gratefully acknowledge the University of Bath’s Research Computing Group (doi.org/10.15125/b6cd-s854) for their support in this work.
@@ A.2 Representation of the MGMC iteration
Multiplying the first sum in (A.6) from the left by $A$, using BA=ISBA=I-S and evaluating the telescoping sum we find
@@
Let $A\in\mathbb{R}^{n\times n}$ be symmetric and invertible. Let the matrices $M,N\in\mathbb{R}^{n\times n}$ be such that A=MNA=M-N and that $M$ is invertible. Then, with $S:=M^{-1}N$,
@@
$\mathscr{M}_{\ell}:=\{(m_{\ell},m_{\ell+1},\dots,m_{L-1})\mid 0\leq m_{\ell^{\prime}}\leq\gamma_{\ell^{\prime}+1}-1\;\text{for all $\ell^{\prime}=\ell,\dots,L$}\}$ (A.9)
@@
$W_{\ell}(m_{\ell},\dots,m_{L})$ $=(S^{\mathrm{post}}_{\ell})^{\nu_{2}}\Bigl(I_{\ell-1}^{\ell}\sum_{m=1}^{\gamma_{\ell}}X_{\ell-1}^{m-1}W_{\ell-1}(\gamma_{\ell}-m,m_{\ell},\dots,m_{L})$ (A.12)
@@
[…]
@@ A.4 Equivalence of Multigrid and MGMC
Suppose (i) holds. Then, from [undefaaaa, Theorem 3.5.1], the spectral radius of $X_{L}$ is less than $1$, so that $\lim_{m\to\infty}X_{L}^{m}=0$. But from (4.26a) and (4.26b) in Lemma 4.10, we have
@@
[…]
@@ A.5 Convergence of distributions
see [undefaaab, (6.32)]. Here, we note that for sufficiently large $m$, the continuity of the determinant together with Theorem 4.12 implies $\mathrm{det}(\Sigma_{m})\neq 0$. The continuity of the trace function and determinant implies $\lim_{m\to\infty}\operatorname{tr}(A_{L}\Sigma_{m})=n_{L}$ and $\lim_{m\to\infty}\frac{\mathrm{det}(A_{L}^{-1})}{\mathrm{det}(\Sigma_{m})}=1$. Hence,
@@
Similarly, $\lim_{m\to\infty}{D_{\text{KL}}\bigl((A_{L}^{-1}f_{L},A_{L}^{-1})\parallel(\text{$\mu_{m}$, $\Sigma_{m}$})\bigr)=0}$ holds, where we again note $\mathrm{det}(\Sigma_{m})\neq 0$ for sufficiently large $m$. In view of the Pinsker’s inequality, the total variation distance converges to 0 as well, which implies the weak convergence. The weak convergence can be checked directly with the convergence of characteristic function. ∎